OpenStudy (anonymous):

If the perimeter of a square is 48 miles, what is its area? 12 m2 144 m2 24 m2 576 m2

OpenStudy (anonymous):

144 m^2

OpenStudy (ajprincess):

All the sides of a square are equal. if one side is x then perimeter=x+x+x+x=48 4x=48 find x. area = x^2

OpenStudy (anonymous):

You find the answer by doing 48/4 to get 12. Then do 12*12 to get 144

OpenStudy (ajprincess):

@Abyss19707 it wil be highly appreciated if u guide the asker towards the answer:)